OpenStudy (anonymous):

christopher orders a 3 topping pizza for 15.25, and a 5 topping pizza for 17.75. Write and solve a system of linear equations to find the price of a plain cheese pizza (no toppings) and the cost of each topping

OpenStudy (anonymous):

A plain cheese pizza = X A plain cheese pizza + 1 topping = X + Y A plain cheese pizza + 3 toppings = ? A plain cheese pizza + 5 toppings = ?

OpenStudy (shane_b):

A plain cheese pizza = x \[x + 3y = 15.25\]\[x + 5y = 17.75\]Just solve for x and y using substitution...